Find the simple interest on 500 borrowed for 3 years at 8% per annum.
How do you do it?

Found 2 solutions by Theo, Alan3354:
Answer by Theo(13342)

simple interest formula is i = p * r * n

i is the interest
p is the principal
r is the interest rate per time period
n is the number of time periods

your time period is in years.

r = .08 per year
n = 3 years.
p = 500

formula of i = p * r * n becomes i = 500 * .08 * 3 which results in i = 120.


Answer by Alan3354(69443)

Find the simple interest on 500 borrowed for 3 years at 8% per annum.
How do you do it?
============
Interest = 500*0.08*3 = $120